Two concentric uniform thin spherical shells of masses 71 kg and 20 kg and radii 26 m and 87 m, respectively. What is the gravitational force on a particle of mass 130 kg located
(a) a distance 13 m from the center of the shells: F1= ____pN
(b) a distance 62 m from the center of the shells: F2= _____pN
(c) a distance 189 m: F3= _____pN

(Use G=6.67 x 10-11 N kg2/m2)
CAREFUL! ANSWER IN PICONEWTONS (pN) (1 pN=10 x -12 N)
Answer #1

a] Here particle is inside both shell where  gravitational field is zero,

So F1 = 0 pN

b] Here the particle is inside lighter shell, so its gravitational force will be zero, Force will be only due to heavier shell,

F2 = GMm1/r^2 = 6.67e-11*130*71/62^2 = 160*10^-12 N = 160 pN

c] The particle is outside both shells, so force will be due to both of them,

F3 = GMm1/r^2 + GMm2/r^2

= 6.67e-11*130*71/189^2 + 6.67e-11*130*20/189^2

= 22.1e-12 N = 22.1 pN